# Umlaut information

### v1

@@ IV.7 Examples
$U(\pazocal{W}_{q})=-\log\min_{0\leq p\leq 1}\Big{(}\exp\left({p\log(1-q)+(1-p)\log q}\right)+\exp\left({p\log q+(1-p)\log(1-q)}\right)\Big{)},$ (187)
@@
$U(\pazocal{W}_{q})=-\log\left(2\exp\left(\frac{1}{2}\log q(1-q)\right)\right)=-\log\sqrt{q(1-q)}-\log 2.$ (190)
@@
$E^{\emptyset}(0^{+},\pazocal{W}_{q})$ $=-\min_{0\leq p\leq 1}\left(p^{2}\log(1-q+q)+2p(1-p)\log(2\sqrt{q(1-q)})+(1-p)\log(q+1-q)\right)$ (191)
=$\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize(i)}}}}{{=}}-2\min_{0\leq p\leq 1}p(1-p)\log(2\sqrt{q(1-q)})$
@@
where in (i) we have used that $-cp(1-p)$ with $c>0$ is minimised when $p=1/2$. Hence,
$E^{\rm NS}(0^{+},\pazocal{W}_{q})=U(\pazocal{W}_{q})=2E^{\emptyset}(0^{+},\pazocal{W}_{q}).$ (192)
@@
$\pazocal{W}_{q}(y|x)=\begin{cases}1-q&y=x\\ q&y=\varepsilon\\ 0&\text{otherwise}\end{cases}$ (193)
@@
$\exp\left({-U(\pazocal{W}_{q})}\right)$ $=\min_{0\leq p\leq 1}q^{p}q^{1-p}=q$ (194)
@@
$E^{\emptyset}(0^{+},\pazocal{W}_{q})=-\min_{0\leq p\leq 1}2p(1-p)\log q=-\frac{1}{2}\log q,$ (196)
@@
$E^{\rm NS}(0^{+},\pazocal{W}_{q})=U(\pazocal{W}_{q})=2E^{\emptyset}(0^{+},\pazocal{W}_{q}).$ (197)
@@
$\pazocal{W}_{H,m,V}(y|x)=\mathcal{G}(m+Hx,V)(y)\qquad\forall\,x\in\mathbb{R}^{n},\,y\in\mathbb{R}^{k}.$ (199)
@@
[…]
@@
[…]
@@
[…]
@@
$=+\frac{1}{2}\operatorname{Tr}[CH^{\intercal}V^{-1}H]\,,$
@@
IV.8 Comparison with channel lautum information
@@
$P_{X^{n}Y^{n}}(x_{1},\dots,x_{n},y_{1},\dots,y_{n})=P_{X}(x_{1},\dots,x_{n})\prod_{i=1}^{n}\pazocal{W}(y_{i}|x_{i}).$ (205)
@@
$L^{\infty}(\pazocal{W})$ $\coloneqq\liminf_{n\to\infty}\frac{1}{n}L(\pazocal{W}^{\times n}).$ (208)
@@
$E^{\emptyset}(0^{+},\pazocal{W})\leq L^{\infty}(\pazocal{W}).$ (210)
@@
$L^{\infty}(\pazocal{W})=\liminf_{n\to\infty}\frac{1}{n}L(\pazocal{W}^{\times n})\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize(i)}}}}{{\geq}}\liminf_{n\to\infty}\frac{1}{n}U(\pazocal{W}^{\times n})\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize(ii)}}}}{{=}}U(\pazocal{W})\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize(iii)}}}}{{=}}E^{\rm NS}(0^{+},\pazocal{W})\geq E^{\emptyset}(0^{+},\pazocal{W}),$ (211)
@@
$L(\pazocal{W}_{q})$ $=\frac{1}{2}\log\frac{1}{4q(1-q)},$ (213)
$L^{\infty}(\pazocal{W}_{q})$ $=\left(\frac{1}{2}-q\right)\log\frac{1-q}{q}.$
@@
$U(\pazocal{W}_{q})=\frac{1}{2}\log\frac{1}{4q(1-q)},$ (214)
@@ Acknowledgements.
We thank Hao-Chung Cheng for enlightening correspondence on low-rate codes. MB acknowledges funding from the European Research Council (ERC Grant Agreement No. 948139) and the Excellence Cluster Matter andLight for Quantum Computing (ML4Q). LL acknowledges financial support from the European Union under the European Research Council (ERC Grant Agreement No. 101165230) and from MIUR (Ministero dell’Istruzione, dell’Università e della Ricerca) through the project ‘Dipartimenti di Eccellenza 2023–2027’ of the ‘Classe di Scienze’ department at the Scuola Normale Superiore. MT is supported by the Ministry of Education through grant T2EP20124-0005 and by the National Research Foundation, Singapore through the National Quantum Office, hosted in A*STAR, under its Centre for Quantum Technologies Funding Initiative (S24Q2d0009). AO acknowledges funding by the European Research Council (ERC Grant Agreement No. 948139).
@@ Appendix A Auxiliary proofs
$E_{0}^{\rm NS}(\pazocal{W})$ $=\lim_{\delta\to 0}\liminf_{n\to\infty}\frac{1}{n}\max_{P_{X^{n}}}\min_{Q_{Y^{n}}}D^{\delta}_{H}(P_{X^{n}}Q_{Y^{n}}\|P_{X^{n}Y^{n}})$ (216)
@@
$\varepsilon^{\rm NS}(M,\pazocal{W})$ $=1-\max_{R,P_{X}}\Bigg{\{}\begin{aligned} &\sum_{x\in\mathcal{X}}\sum_{y\in\mathcal{Y}}\pazocal{W}(y|x)\,R_{xy}\,:\\ &\sum_{x\in\mathcal{X}}R_{xy}\leq\frac{1}{M}\quad\forall y\in\mathcal{Y},\\ &0\leq R_{xy}\leq P_{X}(x)\quad\forall x\in\mathcal{X},y\in\mathcal{Y},\\ &P_{X}\in\mathcal{P}(\mathcal{X})\Bigg{\}}\end{aligned}$ (217)
$=\min_{R^{\prime},P_{X}}\Bigg{\{}\begin{aligned} &\sum_{x\in\mathcal{X}}\sum_{y\in\mathcal{Y}}\pazocal{W}(y|x)\,R^{\prime}_{xy}\,:\\ &\sum_{x\in\mathcal{X}}R^{\prime}_{xy}\geq 1-\frac{1}{M}\quad\forall y\in\mathcal{Y},\\ &0\leq R^{\prime}_{xy}\leq P_{X}(x)\quad\forall x\in\mathcal{X},y\in\mathcal{Y},\\ &P_{X}\in\mathcal{P}(\mathcal{X})\Bigg{\}},\end{aligned}$ (218)
@@
$-\log\varepsilon^{\rm NS}(M,\pazocal{W})$ $=-\log\min_{T,P_{X}}\Bigg{\{}\begin{aligned} &\sum_{x\in\mathcal{X}}\sum_{y\in\mathcal{Y}}\pazocal{W}(y|x)\,P_{X}(x)\,T_{xy}\,:\\ &\sum_{x\in\mathcal{X}}P_{X}(x)\,T_{xy}\geq 1-\frac{1}{M}\quad\forall y\in\mathcal{Y},\\ &0\leq T_{xy}\leq 1\quad\forall x\in\mathcal{X},y\in\mathcal{Y},\\ &P_{X}\in\mathcal{P}(\mathcal{X})\Bigg{\}}\end{aligned}$ (219)
$=-\log\min_{P_{X}\in\mathcal{P}(\mathcal{X})}\min_{T}\Bigg{\{}\begin{aligned} &\sum_{x\in\mathcal{X}}\sum_{y\in\mathcal{Y}}P_{XY}(x,y)\,T_{xy}\,:\\ &\min_{Q_{Y}\in\mathcal{P}(\mathcal{Y})}\sum_{x\in\mathcal{X}}P_{X}(x)\,Q_{Y}(y)\,T_{xy}\geq 1-\frac{1}{M},\\ &0\leq T_{xy}\leq 1\quad\forall x\in\mathcal{X},y\in\mathcal{Y}\Bigg{\}}\end{aligned}$ (220)
$=-\log\min_{P_{X}\in\mathcal{P}(\mathcal{X})}\max_{Q_{Y}\in\mathcal{P}(\mathcal{Y})}\min_{T}\Bigg{\{}\begin{aligned} &\sum_{x\in\mathcal{X}}\sum_{y\in\mathcal{Y}}P_{XY}(x,y)\,T_{xy}\,:\\ &\sum_{x\in\mathcal{X}}P_{X}(x)\,Q_{Y}(y)\,T_{xy}\geq 1-\frac{1}{M},\\ &0\leq T_{xy}\leq 1\quad\forall x\in\mathcal{X},y\in\mathcal{Y}\Bigg{\}}\end{aligned}$ (221)
$=\max_{P_{X}\in\mathcal{P}(\mathcal{X})}\min_{Q_{Y}\in\mathcal{P}(\mathcal{Y})}D^{1/M}_{H}(P_{X}Q_{Y}\|P_{XY}),$ (222)
@@
$E_{0}^{\rm NS}(\pazocal{W})$ $=\lim_{M\to\infty}\liminf_{n\to\infty}-\frac{1}{n}\log\varepsilon^{\rm NS}(M,\pazocal{W}^{\times n})$ (223)

### v2

@@ IV.8 Examples
$\hskip 0.0ptU(\pazocal{W}_{q})=-\log\min_{0\leq p\leq 1}\Big{(}\exp\left({p\log(1-q)+(1-p)\log q}\right)+\exp\left({p\log q+(1-p)\log(1-q)}\right)\Big{)},$ (226)
@@
$\hskip 0.0ptU(\pazocal{W}_{q})=-\log\left(2\exp\left(\frac{1}{2}\log q(1-q)\right)\right)=-\log\sqrt{q(1-q)}-\log 2.$ (229)
@@
$\hskip 0.0ptE^{\emptyset}(0^{+},\pazocal{W}_{q})$ $=-\min_{0\leq p\leq 1}\left(p^{2}\log(1-q+q)+2p(1-p)\log(2\sqrt{q(1-q)})+(1-p)\log(q+1-q)\right)$ (230)
$\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize(i)}}}}{{=}}-2\min_{0\leq p\leq 1}p(1-p)\log(2\sqrt{q(1-q)})$
@@
where in (i) we have used that cp(1p)-cp(1-p) with $c>0$ is minimised when $p=1/2$. Hence,
$\hskip 0.0ptE^{\rm NS}(0^{+},\pazocal{W}_{q})=U(\pazocal{W}_{q})=2E^{\emptyset}(0^{+},\pazocal{W}_{q}).$ (231)
@@
[…]
@@
$\hskip 0.0pt\exp\left({-U(\pazocal{W}_{q})}\right)$ $=\min_{0\leq p\leq 1}q^{p}q^{1-p}=q$ (233)
@@
$\hskip 0.0ptE^{\emptyset}(0^{+},\pazocal{W}_{q})=-\min_{0\leq p\leq 1}2p(1-p)\log q=-\frac{1}{2}\log q,$ (235)
@@
$\hskip 0.0ptE^{\rm NS}(0^{+},\pazocal{W}_{q})=U(\pazocal{W}_{q})=2E^{\emptyset}(0^{+},\pazocal{W}_{q}).$ (236)
@@
$\hskip 0.0pt\pazocal{W}_{H,m,V}(y|x)=\mathcal{G}(m+Hx,V)(y)\qquad\forall\,x\in\mathbb{R}^{n},\,y\in\mathbb{R}^{k}.$ (238)
@@
$\hskip 0.0ptU(\pazocal{W}_{H,m,V})$ $=-\log\min_{p_{X}}\int_{\mathbb{R}^{k}}dy\,\exp\left(\int_{\mathbb{R}^{n}}dx\,p_{X}(x)\log\mathcal{G}(m+Hx,V)(y)\right)$ (239)
$=-\log\min_{p_{X}}\int_{\mathbb{R}^{k}}dy\,\exp\left(-\frac{1}{2}\int_{\mathbb{R}^{n}}dx\,p_{X}(x)(y-m-Hx)^{\intercal}V^{-1}(y-m-Hx)\right)$
@@
$\hskip 0.0pt\int_{\mathbb{R}^{n}}dx$ $\,p_{X}(x)(y-m-Hx)^{\intercal}V^{-1}(y-m-Hx)$ (240)
$=\int_{\mathbb{R}^{n}}dx\,p_{X}(x)\operatorname{Tr}[V^{-1}(y-m-Hx)(y-m-Hx)^{\intercal}]$
@@
$\hskip 0.0ptU(\pazocal{W}_{H,m,V})$ $=-\log\min_{\mu}\frac{1}{\sqrt{(2\pi)^{n}\det V}}\int_{\mathbb{R}^{k}}dy\,\exp\left(-\frac{1}{2}(y-m-H\mu)^{\intercal}V^{-1}(y-m-H\mu)\right)$ (241)
@@
$=\frac{1}{2}\operatorname{Tr}[CH^{\intercal}V^{-1}H]\,$,
@@
IV.9 Comparison with channel lautum information
@@
$\hskip 0.0ptP_{X^{n}Y^{n}}(x_{1},\dots,x_{n},y_{1},\dots,y_{n})=P_{X}(x_{1},\dots,x_{n})\prod_{i=1}^{n}\pazocal{W}(y_{i}|x_{i}).$ (244)
@@
$\hskip 0.0ptL^{\infty}(\pazocal{W})$ $\coloneqq\liminf_{n\to\infty}\frac{1}{n}L(\pazocal{W}^{\times n}).$ (247)
@@
$\hskip 0.0ptE^{\emptyset}(0^{+},\pazocal{W})\leq L^{\infty}(\pazocal{W}).$ (249)
@@
$\hskip 0.0ptL^{\infty}(\pazocal{W})=\liminf_{n\to\infty}\frac{1}{n}L(\pazocal{W}^{\times n})\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize(i)}}}}{{\geq}}\liminf_{n\to\infty}\frac{1}{n}U(\pazocal{W}^{\times n})\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize(ii)}}}}{{=}}U(\pazocal{W})\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize(iii)}}}}{{=}}E^{\rm NS}(0^{+},\pazocal{W})\geq E^{\emptyset}(0^{+},\pazocal{W}),$ (250)
@@
$\hskip 0.0ptL(\pazocal{W}_{q})$ $=\frac{1}{2}\log\frac{1}{4q(1-q)},L^{\infty}(\pazocal{W}_{q})$ $=\left(\frac{1}{2}-q\right)\log\frac{1-q}{q}.$ (252)
@@
$\hskip 0.0ptU(\pazocal{W}_{q})=\frac{1}{2}\log\frac{1}{4q(1-q)},$ (253)
@@ Acknowledgements.
We thank Hao-Chung Cheng for enlightening correspondence on low-rate codes. MB and AO acknowledge funding from the European Research Council (ERC Grant Agreement No. 948139) and the Excellence Cluster Matter and Light for Quantum Computing (ML4Q). LL acknowledges financial support from the European Union under the European Research Council (ERC Grant Agreement No. 101165230) and from MIUR (Ministero dell’Istruzione, dell’Università e della Ricerca) through the project ‘Dipartimenti di Eccellenza 2023–2027’ of the ‘Classe di Scienze’ department at the Scuola Normale Superiore. MT is supported by the Ministry of Education through grant T2EP20124-0005 and by the National Research Foundation, Singapore through the National Quantum Office, hosted in A*STAR, under its Centre for Quantum Technologies Funding Initiative (S24Q2d0009).
@@ Appendix A Auxiliary proofs
$\hskip 0.0ptE_{0}^{\rm NS}(\pazocal{W})$ $=\lim_{\delta\to 0}\liminf_{n\to\infty}\frac{1}{n}\max_{P_{X^{n}}}\min_{Q_{Y^{n}}}D^{\delta}_{H}(P_{X^{n}}Q_{Y^{n}}\|P_{X^{n}Y^{n}})$ (255)
@@
$\hskip 0.0pt\varepsilon^{\rm NS}(M,\pazocal{W})$ $=1-\max_{\begin{subarray}{c}R\\ P_{X}\in\mathcal{P}(\mathcal{X})\end{subarray}}\Bigg{\{}\begin{aligned} &\sum_{x\in\mathcal{X}}\sum_{y\in\mathcal{Y}}\pazocal{W}(y|x)\,R_{xy}\,:\\ &\sum_{x\in\mathcal{X}}R_{xy}\leq\frac{1}{M}\quad\forall y\in\mathcal{Y},\quad 0\leq R_{xy}\leq P_{X}(x)\quad\forall x\in\mathcal{X},y\in\mathcal{Y}\Bigg{\}}\end{aligned}$ (256)
$=\min_{\begin{subarray}{c}R^{\prime}\\ P_{X}\in\mathcal{P}(\mathcal{X})\end{subarray}}\Bigg{\{}\begin{aligned} &\sum_{x\in\mathcal{X}}\sum_{y\in\mathcal{Y}}\pazocal{W}(y|x)\,R^{\prime}_{xy}\,:\\ &\sum_{x\in\mathcal{X}}R^{\prime}_{xy}\geq 1-\frac{1}{M}\quad\forall y\in\mathcal{Y},\quad 0\leq R^{\prime}_{xy}\leq P_{X}(x)\quad\forall x\in\mathcal{X},y\in\mathcal{Y}\Bigg{\}},\end{aligned}$
@@
$\hskip 0.0pt-\log$ $\varepsilon^{\rm NS}(M,\pazocal{W})$ (257)
$=-\log\min_{\begin{subarray}{c}T\\ P_{X}\in\mathcal{P}(\mathcal{X})\end{subarray}}\Bigg{\{}\begin{aligned} &\sum_{x\in\mathcal{X}}\sum_{y\in\mathcal{Y}}\pazocal{W}(y|x)\,P_{X}(x)\,T_{xy}\,:\\ &\sum_{x\in\mathcal{X}}P_{X}(x)\,T_{xy}\geq 1-\frac{1}{M}\quad\forall y\in\mathcal{Y},\quad 0\leq T_{xy}\leq 1\quad\forall x\in\mathcal{X},y\in\mathcal{Y}\Bigg{\}}\end{aligned}$
$=-\log\min_{P_{X}\in\mathcal{P}(\mathcal{X})}\min_{T}\Bigg{\{}\begin{aligned} &\sum_{x\in\mathcal{X}}\sum_{y\in\mathcal{Y}}P_{XY}(x,y)\,T_{xy}\,:\\ &\min_{Q_{Y}\in\mathcal{P}(\mathcal{Y})}\sum_{x\in\mathcal{X}}P_{X}(x)\,Q_{Y}(y)\,T_{xy}\geq 1-\frac{1}{M},\quad 0\leq T_{xy}\leq 1\quad\forall x\in\mathcal{X},y\in\mathcal{Y}\Bigg{\}}\end{aligned}$
$=-\log\min_{P_{X}\in\mathcal{P}(\mathcal{X})}\max_{Q_{Y}\in\mathcal{P}(\mathcal{Y})}\min_{T}\Bigg{\{}\begin{aligned} &\sum_{x\in\mathcal{X}}\sum_{y\in\mathcal{Y}}P_{XY}(x,y)\,T_{xy}\,:\\ &\sum_{x\in\mathcal{X}}P_{X}(x)\,Q_{Y}(y)\,T_{xy}\geq 1-\frac{1}{M},\quad 0\leq T_{xy}\leq 1\quad\forall x\in\mathcal{X},y\in\mathcal{Y}\Bigg{\}}\end{aligned}$
$=\max_{P_{X}\in\mathcal{P}(\mathcal{X})}\min_{Q_{Y}\in\mathcal{P}(\mathcal{Y})}D^{1/M}_{H}(P_{X}Q_{Y}\|P_{XY}),$
@@
$\hskip 0.0ptE_{0}^{\rm NS}(\pazocal{W})$ $=\lim_{M\to\infty}\liminf_{n\to\infty}-\frac{1}{n}\log\varepsilon^{\rm NS}(M,\pazocal{W}^{\times n})$ (258)